OpenStudy (anonymous):

Let f(x)=2xsquared+1 and g(x)=x-1/x Answer choices: -6 33/32 32/33 -2

OpenStudy (anonymous):

FIND (g o f)(-4)

OpenStudy (accessdenied):

\[ g(x) = \frac{x-1}{x} \] ?

OpenStudy (anonymous):

yea

OpenStudy (accessdenied):

(g o f)(-4) = g( f(-4) ) f(-4) = 2(-4)^2 + 1 = 2(16) + 1 = 32 + 1 = 33 ( insert back into g(x) ) g(33) = (33 - 1)/33 = 32/33

OpenStudy (anonymous):

yea same